# Scalable Computations for Generalized Mixed Effects Models with Crossed Random Effects Using Krylov Subspace Methods

### v1

@@ 4.1 SSOR preconditioner
$1-\max((D_{1}^{-1})_{ii})\max((D_{2}^{-1})_{ii})\max((Z^{T}WZ)_{ii})^{2}\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\min((D_{1}^{-1})_{ii})\min((D_{2}^{-1})_{ii})\min((Z^{T}WZ)_{ii})^{2}.$ (24)
A proof of Theorem 4.2 can be found in Appendix A.1. Theorem 4.2 implies the following corollary for Gaussian likelihoods.
Corollary 4.1.
@@
$\lambda_{m_{1}}^{\text{SSOR}}=\dots=\lambda_{1}^{\text{SSOR}}=\lambda_{\max}^{\text{SSOR}}=1,$
@@
$1-\left(\frac{\max(\sigma^{2}_{1},\sigma^{2}_{2})d^{\max}}{\sigma^{2}+\max(\sigma^{2}_{1},\sigma^{2}_{2})d^{max}}\right)^{2}\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\left(\frac{\min(\sigma^{2}_{1},\sigma^{2}_{2})d^{\min}}{\sigma^{2}+\min(\sigma^{2}_{1},\sigma^{2}_{2})d^{\min}}\right)^{2},$ (25)
where $d^{\max}=\max(Z^{T}Z)$ and $d^{\min}=\min(Z^{T}Z)$ are the largest and smallest number of repeated measurements for the two grouped random effects, respectively.
If, in addition, the design of the random effects is balanced such that every random effect realization occurs $d$ times, i.e., $d=\frac{n}{m_{k}}=\sum_{i=1}^{n}(Z_{k})_{ij}$ for all $1\leq j\leq m_{k}$ and $k=1,2$, the following holds true:
@@
$\lambda_{\min}^{\text{SSOR}}$ $=1-\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{1}d}+1}\right)\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{2}d}+1}\right)$ (26)
Proof of Corollary 4.1.
The statement in (25) follows from (23). Similarly, (26) follows from (24) by plugging in $D_{k}=\text{diag}(1/\sigma^{2}_{k}+d_{k}/\sigma^{2})$, $k=1,2$. ∎
[…]
@@
If $K=2$, the likelihood is Gaussian, and $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a bipartite, biregular random graph with uniform distribution over all bipartite, biregular random graphs, the following hold:
$\lambda_{m-1}^{\text{SSOR}}\geq 1-\frac{(\sqrt{d_{1}-1}+\sqrt{d_{2}-1})^{2}}{(\sigma^{2}/\sigma_{1}^{2}+d_{1})(\sigma^{2}/\sigma_{2}^{2}+d_{2})}-\epsilon_{m}$ (27)
and
@@
A proof of Theorem 4.3 can be found in Appendix A.1. Pandolfi et al. (2024, Theorem 4) derive a similar result for the diagonal preconditioner. However, the upper bound for the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}$ of the SSOR-preconditioned matrix given by $1/(1-(\frac{1}{d_{1}}+\frac{1}{d_{2}}+\frac{2}{\sqrt{d_{1}d_{2}}})-\epsilon_{m})$ decreases faster than the corresponding upper bound $(1+\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}+\epsilon_{m})/(1-(\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}})-\epsilon_{m})$ for the diagonal preconditioner derived in Pandolfi et al. (2024, Theorem 4); see Section 4.3 for more details. Note that the random effects model used in Pandolfi et al. (2024) contains a random intercept term, which we do not include, and they derive their result for $\kappa_{m-2,3}^{\text{Diag}}$ instead of $\kappa_{m-1,1}^{\text{Diag}}$. However, in Theorem A.1 in Appendix A.2, we show that the same result holds true for $\kappa_{m-1,1}^{\text{Diag}}$ for the model without a random intercept used in this article.
4.2 Diagonal preconditioner and no preconditioner
We have the following results for the diagonal preconditioner and the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WT$. We comment on these results and compare them in Section 4.3.
Theorem 4.4.
For the diagonal preconditioner $P_{\text{Diag}}=D$, it holds that
$\frac{1}{\max(\Sigma_{ii}D_{ii})}\leq\lambda_{m+1-k}^{\text{Diag}}\leq\frac{1}{\min(\Sigma_{ii}D_{ii})},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1.$ (30)
If the likelihood is Gaussian, we can additionally obtain the following result.
Theorem 4.5.
If the likelihood is Gaussian, it holds that
$1+\frac{K-1}{\frac{\max(\Sigma^{-1})\sigma^{2}}{d^{\min}}+1}\leq\lambda_{\max}^{\text{Diag}}\leq 1+\frac{K-1}{\frac{\min(\Sigma^{-1})\sigma^{2}}{d^{\max}}+1},$ (31)
$\frac{1}{\frac{\max(\Sigma)}{\sigma^{2}}d^{\max}+1}\leq\lambda_{m+1-k}^{\text{Diag}}\leq\frac{1}{\frac{\min(\Sigma)}{\sigma^{2}}d^{\min}+1},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1,$ (32)
where $d^{\max}=\max(Z^{T}Z)$ and $d^{\min}=\min(Z^{T}Z)$.
If, in addition, $K=2$ and the design of the random effects is balanced such that every random effect realization occurs $d_{k}$ times, i.e., $d_{k}=\frac{n}{m_{k}}=\sum_{i=1}^{n}(Z_{k})_{ij}$ for all $1\leq j\leq m_{k}$ and $k=1,2$, it holds that
Theorem 4.6.
The following hold true for the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WZ$:
$\frac{1}{\max(\Sigma_{ii})}\leq\lambda_{m+1-k}^{\text{none}}\leq\frac{1}{\min(\Sigma_{ii})},\leavevmode\nobreak\ \leavevmode\nobreak\ k=1,\dots,K-1.$ (36)
Proofs of Theorems 4.4, 4.5, and 4.6 can be found in Appendix A.1.
4.3 Comparison of preconditioners
The above results show that the largest and smallest eigenvalues of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for the SSOR, the diagonal, and no preconditioner depend only on the number of repeated observations per random effect, denoted by $d_{k}=\frac{n}{m_{k}}$ when having a balanced design, and not directly on neither the sample size $n$ nor the dimension $m$ of $\Sigma^{-1}+Z^{T}WZ$. In the following, we analyze and compare the extremal eigenvalues and condition numbers of the preconditioned matrices in more detail.
Theorem 4.7 below demonstrates that, for $K=2$, a Gaussian likelihood, and a balanced design with $d=m_{1}/m_{2}$ repeated observations per random effect, the condition numbers $\kappa=\lambda_{1}/\lambda_{m}$ for the SSOR and diagonal preconditioners and also when not applying any preconditioner grow asymptotically linearly in $d$. In detail, Theorem 4.7 shows that the condition number and the difference of the largest and smallest eigenvalue of the preconditioned matrix $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$, which matter for SLQ, are smaller for the SSOR preconditioner compared to the diagonal preconditioner. Furthermore, the theorem shows that the condition number $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ of the SSOR-preconditioned matrix has a linear asymptote in $d$ with a slope that is four times smaller than the corresponding slope of the diagonal preconditioner. In other words, the condition number of the diagonal-preconditioned matrix grows four times faster in $d$ than the one of the SSOR-preconditioned matrix. Based on this, we expect the SSOR preconditioner to result in more accurate SLQ-approximated log-determinants compared to the diagonal preconditioner. This is confirmed empirically in Section 5.
Furthermore, we can obtain an explicit expression for the condition number of the non-preconditioned matrix $\Sigma^{-1}+Z^{T}WZ$ if the two random effects have equal variances, $\sigma^{2}_{1}=\sigma^{2}_{2}$ as the corresponding inequalities in Theorem 4.7 then become equalities. In this case, the condition number of the non-preconditioned and the diagonal-preconditioned matrix are equal and given by $\frac{2\sigma_{1}^{2}}{\sigma^{2}}d+1$, whereas the condition number of the SSOR-preconditioned matrix asymptotically grows with a rate of $\frac{\sigma^{2}_{1}}{2\sigma^{2}}$ in $d$. For the diagonal preconditioner, this can be seen by noting that $\lambda_{\max}^{\text{Diag}}/\lambda_{\min}^{\text{Diag}}=((\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1)+1)/((\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1)-1)=\frac{2\sigma_{1}^{2}}{\sigma^{2}}d+1$ due to (49) in the proof of Theorem 4.7 in Appendix A.1. We conjecture that similar results also hold for the non-preconditioned matrix and the diagonal preconditioner for the general case $\sigma^{2}_{1}\neq\sigma^{2}_{2}$.
Theorem 4.7.
If $K=2$, the likelihood is Gaussian, and the random effects have a balanced design with $d=d_{1}=\frac{n}{m_{1}}=d_{2}=\frac{n}{m_{2}}$ repeated observations per random effect, the following hold true:
$\lambda_{\max}^{\text{SSOR}}<\lambda_{\max}^{\text{Diag}},\leavevmode\nobreak\ \leavevmode\nobreak\ \lambda_{\min}^{\text{SSOR}}>\lambda_{\min}^{\text{Diag}},$ (37)
$\lim_{d\rightarrow\infty}\left(\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}-\left(\frac{\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1\right)\right)=0,$ (38)
@@ 4.3 Comparison of preconditioners
$\lim_{d\rightarrow\infty}\left(\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}-\left(\frac{4\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1\right)\right)=0,$ (39)
$\frac{2\min(\sigma^{2}_{1},\sigma^{2}_{2})}{\sigma^{2}}d+\frac{\min(\sigma^{2}_{1},\sigma^{2}_{2})}{\max(\sigma^{2}_{1},\sigma^{2}_{2})}\leq\frac{\lambda_{\max}^{\text{none}}}{\lambda_{\min}^{\text{none}}}\leq\frac{2\max(\sigma^{2}_{1},\sigma^{2}_{2})}{\sigma^{2}}d+\frac{\max(\sigma^{2}_{1},\sigma^{2}_{2})}{\min(\sigma^{2}_{1},\sigma^{2}_{2})},$ (40)
@@
We can additionally obtain upper bounds for the effective condition numbers $\kappa_{m-1,1}^{\text{SSOR}}$ and $\kappa_{m-1,2}^{\text{Diag}}$ for the SSOR and diagonal preconditioner, respectively, and a Gaussian likelihood when assuming that $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a bipartite, biregular random graph. Theorem 4.8 below shows that the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}$ of the SSOR-preconditioned matrix essentially decreases with a rate of $\frac{4}{d}$, up to higher-order terms. In contrast, the effective condition number $\kappa_{m-1,2}^{\text{Diag}}$ of the diagonal preconditioner decreases with a slower rate of $\frac{4}{\sqrt{d}}$, again up to higher-order terms. Note that the assumption $m_{1}=m_{2}$ in Theorem 4.8 allows for simpler results but can be easily relaxed. Based on these results, we expect the SSOR preconditioner to lead to faster convergence of the CG method compared to the diagonal preconditioner. A proof of Theorem 4.8 can be found in Appendix A.2.
Theorem 4.8.
@@
$\kappa_{m-1,2}^{\text{Diag}}=\frac{\lambda_{2}^{\text{Diag}}}{\lambda_{m-1}^{\text{Diag}}}$ $\leq\frac{1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}}{1-\frac{2}{\sqrt{d}}-\epsilon_{m}^{\prime}}=1+\frac{4}{\sqrt{d}}+2\epsilon_{m}^{\prime}+O\left(\frac{1}{d},\epsilon_{m}^{\prime 2},\frac{\epsilon_{m}^{\prime}}{\sqrt{d}}\right)$ (42)
asymptotically almost surely with $\epsilon_{m}\rightarrow 0$ and $\epsilon_{m}^{\prime}\rightarrow 0$ as $m\rightarrow\infty$.

### v2

@@ 4.1 Extremal eigenvalues and convergence of the SLQ method
$1-\frac{\max((Z^{T}WZ)_{ii})^{2}}{\min((D_{1})_{ii})\min((D_{2})_{ii})}$ $\leq\lambda_{\min}^{\text{SSOR}}\leq 1-\frac{\min((Z^{T}WZ)_{ii})^{2}}{\max((D_{1})_{ii})\max((D_{2})_{ii})},$ (23)
where $D_{1}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})\in\mathbb{R}^{m_{1}\times m_{1}}$ and $D_{2}=\text{diag}((\Sigma^{-1}+Z^{T}WZ)_{ii})\in\mathbb{R}^{(m-m_{1})\times(m-m_{1})}$.
Lemma 4.2.
For the diagonal preconditioner $P_{\text{Diag}}=D$, it holds that
$\frac{1}{1+\max(\Sigma_{ii}(Z^{T}WZ)_{ii})}$ $\leq\lambda_{m+1-k}^{\text{Diag}}\leq\frac{1}{1+\min(\Sigma_{ii}(Z^{T}WZ)_{ii})},\penalty 10000\ \penalty 10000\ k=1,\dots,K-1.$ (25)
Moreover, the following hold true for the unpreconditioned matrix $\Sigma^{-1}+Z^{T}WZ$:
$\frac{1}{\max(\Sigma_{ii})}$ $\leq\lambda_{m+1-k}^{\text{none}}\leq\frac{1}{\min(\Sigma_{ii})},\penalty 10000\ \penalty 10000\ k=1,\dots,K-1.$ (27)
Theorem 4.2.
@@
$\frac{1}{1-\left(\frac{\sigma^{2}_{\min}d_{\min}}{\sigma^{2}+\sigma^{2}_{\min}d_{\min}}\right)^{2}}$ $\leq\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}\leq\frac{1}{1-\left(\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}+\sigma^{2}_{\max}d_{max}}\right)^{2}}=\frac{1}{2}\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}}+\frac{3}{4}+o(1)\penalty 10000\ \penalty 10000\ (d_{\max}\to\infty),$ (28)
$2\frac{\sigma^{2}_{\min}d_{\min}}{\sigma^{2}}+1$ $\leq\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}\leq 2\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}}+1,$ (29)
$2\frac{\sigma^{2}_{\min}d_{\min}}{\sigma^{2}}+\frac{\sigma^{2}_{\min}}{\sigma^{2}_{\max}}$ $\leq\frac{\lambda_{\max}^{\text{none}}}{\lambda_{\min}^{\text{none}}}\leq 2\frac{\sigma^{2}_{\max}d_{\max}}{\sigma^{2}}+\frac{\sigma^{2}_{\max}}{\sigma^{2}_{\min}}.$ (30)
If, in addition, the design of the random effects is balanced, the following hold:
@@
$\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}$ $=\frac{1}{1-\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{1}d}+1}\right)\left(\frac{1}{\frac{\sigma^{2}}{\sigma^{2}_{2}d}+1}\right)}=\frac{\sigma_{1}^{2}\sigma_{2}^{2}}{\sigma^{2}(\sigma_{1}^{2}+\sigma_{2}^{2})}\,d+1+o(1)\penalty 10000\ \penalty 10000\ (d\to\infty),$ (31)
$\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}$ $=\frac{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}+1}{\sqrt{\left(\frac{\sigma^{2}}{\sigma_{1}^{2}d}+1\right)\left(\frac{\sigma^{2}}{\sigma_{2}^{2}d}+1\right)}-1}=\frac{4\sigma^{2}_{1}\sigma^{2}_{2}}{\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2})}d+1+o(1)\penalty 10000\ \penalty 10000\ (d\to\infty).$ (32)
A proof of Theorem 4.2 can be found in Appendix A.1. This theorem has the following implications. First, Theorem 4.2 shows that the condition numbers $\lambda_{\max}^{\text{P}}/\lambda_{\min}^{\text{P}}$ of the SSOR and diagonal preconditioners as well as the unpreconditioned matrix do not directly depend on neither the sample size $n$ nor the dimension of the random effects $m$ but rather only on (i) the number of occurrences per random effect $(Z^{T}Z)_{ii}$ ($d=n/m_{1}=n/m_{2}$ when having a balanced design) and (ii) the signal-to-noise ratios $\sigma^{2}_{k}/\sigma^{2}$ (and $\sigma^{2}_{\max}/\sigma^{2}_{\min}$ when no preconditioning is used). Specifically, we find that the condition numbers are increasing in the number of repeated occurrences per random effect and the signal-to-noise ratios. Concerning the former, the theorem shows that the condition numbers $\lambda_{\max}^{\text{P}}/\lambda_{\min}^{\text{P}}$ of the SSOR and diagonal preconditioners both grow asymptotically linearly in $d$ when having a balanced random effects design. However, $\lambda_{\max}^{\text{Diag}}/\lambda_{\min}^{\text{Diag}}$ asymptotically grows four times faster in $d$ than $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ ; see (31) and (32). Regarding the signal-to-noise ratios, similar arguments as in the derivations of the asymptotic results in (31) and (32) can be made that $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ and $\lambda_{\max}^{\text{Diag}}/\lambda_{\min}^{\text{Diag}}$ both have linear asymptotes in the “average” signal-to-noise ratio $2\sigma^{2}_{1}\sigma^{2}_{2}/(\sigma^{2}(\sigma^{2}_{1}+\sigma^{2}_{2}))$. Again, $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ grows four times slower in this signal-to-noise ratio compared to $\lambda_{\max}^{\text{Diag}}/\lambda_{\min}^{\text{Diag}}$. Analogous arguments concerning $d_{\max}$ and the signal-to-noise ratio can be made for the upper bounds of the condition numbers for general, non-balanced designs.
Moreover, Theorem 4.2 implies
$\lambda_{\max}^{\text{SSOR}}<\lambda_{\max}^{\text{Diag}}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \lambda_{\min}^{\text{SSOR}}>\lambda_{\min}^{\text{Diag}}.$
This holds for all random effects designs and parameter choices. For the balanced design, we can additionally obtain an explicit expression for the condition number of the unpreconditioned matrix $\Sigma^{-1}+Z^{T}WZ$ if the two random effects have equal variances since the inequalities in (30) then become equalities. In this special case, the condition numbers of the unpreconditioned and the diagonal-preconditioned matrix are equal and given by $2\sigma_{1}^{2}d/\sigma^{2}+1$.
Next, in Theorem 4.3, we derive upper bounds for the condition numbers of $P^{-1/2}(\Sigma^{-1}+Z^{T}WZ)P^{-T/2}$ for a Bernoulli likelihood with logit and probit links used in binary classification. Similarly as for a Gaussian likelihood, the upper bound for the condition number $\lambda_{\max}^{\text{SSOR}}/\lambda_{\min}^{\text{SSOR}}$ of the SSOR preconditioner given in (33) is smaller and grows four times slower in both (i) the maximal number of repeated random effects occurrences $d_{\max}$ and (ii) the maximal random effects variance $\sigma^{2}_{\max}$ compared to the diagonal preconditioner and the unpreconditioned matrix given in (34) and (35), respectively. A proof of Theorem 4.3 can be found in Appendix A.1.
@@
If $K=2$ and the likelihood is Bernoulli with a logit or probit link, it holds that
$\frac{\lambda_{\max}^{\text{SSOR}}}{\lambda_{\min}^{\text{SSOR}}}$ $\leq\frac{1}{1-\left(\frac{\sigma^{2}_{\max}d_{\max}C}{1+\sigma^{2}_{\max}d_{\max}C}\right)^{2}}=\frac{1}{2}\sigma^{2}_{\max}d_{\max}C+\frac{3}{4}+o(1)\penalty 10000\ \penalty 10000\ (d_{\max}\to\infty),$ (33)
$\frac{\lambda_{\max}^{\text{Diag}}}{\lambda_{\min}^{\text{Diag}}}$ $\leq 2\sigma^{2}_{\max}d_{\max}C+1,$ (34)
$\frac{\lambda_{\max}^{\text{none}}}{\lambda_{\min}^{\text{none}}}$ $\leq 2\sigma^{2}_{\max}d_{\max}C+\frac{\sigma^{2}_{\max}}{\sigma^{2}_{\min}},$ (35)
where $C$ equals $0.25$ and $1$ for a logit and a probit link, respectively.
In summary, the results in Theorems 4.1, 4.2, and 4.3, imply that the SSOR preconditioner is expected to yield more accurate SLQ-approximated log-determinants compared to applying no preconditioning and to the diagonal preconditioner for both Gaussian and non-Gaussian likelihoods. These findings are empirically confirmed in the experiments in Section 5.
4.2 Effective condition numbers and convergence of the CG method
Under additional assumptions on the random effects design, we derive in Theorem 4.4 a stochastic upper bound for the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}=\lambda_{1}^{\text{SSOR}}/\lambda_{m-1}^{\text{SSOR}}$ of the SSOR preconditioner obtained after removing the smallest eigenvalue. Note that the assumptions of Theorem 4.4, which are the same as in pandolfi2024conjugate, imply that the random effects design is balanced in the sense that every random effect occurs $d_{k}$ times for $k=1,2$, $d_{k}=n/m_{k}=\sum_{i=1}^{n}(Z_{k})_{ij},\penalty 10000\ \forall\penalty 10000\ 1\leq j\leq m_{k}$, and that all pairs of elements of $b_{1}\in\mathbb{R}^{m_{1}}$ and $b_{2}\in\mathbb{R}^{m_{2}}$ co-occur at most once in all observations, i.e., $Z_{2}^{T}Z_{1}\in\{0,1\}^{m_{2}\times m_{1}}$. This theorem shows that $\kappa_{m-1,1}^{\text{SSOR}}$ decreases fast to one as the number of occurrences per random effect $d_{k}$ increases. Consequently, the preconditioned CG method is expected to converge faster for larger $d_{k}$; see the discussion at the beginning of Section 4.
Theorem 4.4.
If $K=2$, the likelihood is Gaussian, and $\begin{pmatrix}0&Z_{1}^{T}Z_{2}\\ Z_{2}^{T}Z_{1}&0\end{pmatrix}\in\{0,1\}^{m\times m}$ is an adjacency matrix of a random bipartite, biregular graph with uniform distribution over all bipartite, biregular graphs, the following holds:
@@ 4.2 Effective condition numbers and convergence of the CG method
A proof of Theorem 4.4 can be found in Appendix A.1. pandolfi2024conjugate derive a similar result for the diagonal preconditioner. However, the upper bound for the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}$ of the SSOR-preconditioned matrix given in (36) decreases faster than the corresponding upper bound for the diagonal preconditioner which is given by
@@
$\kappa_{m-1,2}^{\text{Diag}}=\frac{\lambda_{2}^{\text{Diag}}}{\lambda_{m-1}^{\text{Diag}}}\leq\frac{1+\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}+\epsilon_{m}}{1-\left(\frac{1}{\sqrt{d_{1}}}+\frac{1}{\sqrt{d_{2}}}\right)-\epsilon_{m}}.$ (37)
@@
To make this explicit, for simplicity we consider the case where $m_{1}=m_{2}$ in Corollary 4.1. This corollary shows that the effective condition number $\kappa_{m-1,1}^{\text{SSOR}}$ of the SSOR-preconditioned matrix essentially decreases with a rate of $4/d$, $d=n/m_{1}=n/m_{2}$, up to higher-order terms. In contrast, the effective condition number $\kappa_{m-1,2}^{\text{Diag}}$ of the diagonal preconditioner decreases with a slower rate of $4/\sqrt{d}$, again up to higher-order terms. Based on these results, we expect the SSOR preconditioner to lead to faster convergence for the CG method compared to the diagonal preconditioner. A proof of Corollary 4.1 can be found in Appendix A.1. Note that the random effects model used in pandolfi2024conjugate contains a random intercept term, which we do not include, and they derive their result for $\kappa_{m-2,3}^{\text{Diag}}$ instead of $\kappa_{m-1,2}^{\text{Diag}}$. However, in Theorem A.2 in Appendix A.1, we show that the result in (37) analogously holds true for the model without a random intercept used in this article.
Corollary 4.1.
@@
$\kappa_{m-1,2}^{\text{Diag}}=\frac{\lambda_{2}^{\text{Diag}}}{\lambda_{m-1}^{\text{Diag}}}$ $\leq\frac{1+\frac{2}{\sqrt{d}}+\epsilon_{m}^{\prime}}{1-\frac{2}{\sqrt{d}}-\epsilon_{m}^{\prime}}=1+\frac{4}{\sqrt{d}}+\epsilon_{m}^{\prime}+O\left(\frac{1}{d},\epsilon_{m}^{\prime 2},\frac{\epsilon_{m}^{\prime}}{\sqrt{d}}\right)$ (39)
asymptotically almost surely with $\epsilon_{m}\rightarrow 0$ and $\epsilon_{m}^{\prime}\rightarrow 0$ as $m\rightarrow\infty$, where $d=n/m_{1}=n/m_{2}$.